what is the quation of a line with a slope of two thirds which contains pont (0,-4)
Answer by stanbon(75887)

What is the equation of a line with a slope
of two thirds which contains point (0,-4).
---------------------
Form: y = mx + b
m = 2/3
b = -4
-------
y = (2/3)x-4
===============
Cheers,
Stan H.
===============